# STT-GS: Sample-Then-Transmit Edge Gaussian Splatting with Joint Client Selection and Power Control

### v3

@@ IV-A Feature Domain Clustering for Efficient Sampling
Specifically, we consider the Hue Saturation Value (HSV) feature representation [kuehni2008color], where H is hue angle, S is saturation intensity, and V is value brightness. The feature extraction process is thus
@@ V EGS Full Transmission With
Problem $\mathsf{P}_{2}$ involves nonlinear coupling among binary variables $\mathbf{x}$ and continuous variables $\mathbf{p}$. Furthermore, even if we relax $\mathbf{x}$ into a continuous variable, the resultant problem is still nonconvex due to potential multi-user interference. Hence, existing optimization solvers such as Mosek are not applicable. To tackle the challenge, in the following, we will propose a PAMM algorithm that solve $\mathsf{P}_{2}$ efficiently. It consists of an outer penalty alternating minimization (PAM) iteration and an inner majorization minimization ($M$) iteration (thus we name it PAMM). Below we present PAM and $M$ in detail.
@@ V-A Penalty Alternating Minimization
To tackle the discontinuity in constraint (12d), we first relax the binary constraint $x_{k}\in\{0,1\}$ into an affine constraint $x_{k}\in[0,1]$, $\forall k$. However, the relaxation is generally not tight and the solution to the relaxed problem could be $0<x_{k}<1$. To promote a binary solution for the relaxed variable $\{x_{k}\}$, we augment the objective function with a penalty term as in [rinaldi2009new]. Here, we adopt the following penalty function [lucidi2010exact]:
@@
According to [lucidi2010exact, Proposition 1], with the penalty term in (24), there exists an upper bound $\bar{\beta}>0$ such that for any $\beta\in[0,\bar{\beta}]$, $\mathrm{P}_{2}$ and $\mathrm{P_{3}}$ are equivalent with a proper choice of $\beta$ [lucidi2010exact].
Now, variables $\{\mathbf{x},\bm{\xi}\}$ and $\mathbf{p}$ are independent in all constraints. Consequently, to solve $\mathrm{P}_{3}$, we can adopt alternating minimization that solves $\{\mathbf{x},\bm{\xi}\}$’s subproblem and $\mathbf{p}$’s subproblem iteratively. This procedure is guaranteed to converge to a local minimum solution of $\mathsf{P}_{3}$ according to [abeck].
@@ V-B Majorization Minimization
In the iterative procedure (28), $\mathrm{P}_{3b}^{[i]}$ is a linearly constrained quadratic optimization problem and can be optimally solved by off-the-shelf software (e.g., CVXPY). However, $\mathrm{P}_{3a}^{[i]}$ is a nonconvex optimization problem, since the function $\varphi_{1}$ in (34a) and the function $\Phi_{k}$ in (34b) are nonconvex. To tackle these functions, we propose to leverage the framework of $M$, which constructs a sequence of upper bounds $\{\widehat{\varphi}_{1}\}$ on $\varphi_{1}(\mathbf{x})$ and replaces $\varphi_{1}(\mathbf{x})$ in $\rm{P}3$ with $\{\widehat{\varphi}_{1}\}$ to obtain the surrogate problems. Similarly, $M$ would also construct upper bounds $\{\widehat{\Phi}_{k}\}$ on $\Phi_{k}$ and replace $\Phi_{k}$ in $\rm{P}3$ with $\{\widehat{\Phi}_{k}\}$. More specifically, given any feasible solution $\{\mathbf{x}^{\star},\bm{\xi}^{\star}\}$ to $\mathrm{P}3$, we define surrogate functions
@@
Based on part (ii) of Proposition 1, the problem $\mathrm{P}_{3a}^{[i]}[n+1]$ is convex and can be solved by off-the-shelf software packages (e.g., Mosek) for convex programming. Denote its optimal solution as $\{\mathbf{x}^{*},\bm{\xi}\}$. Then we set $\mathbf{x}^{[n+1]}=\mathbf{x}^{*}$ and $\{\bm{\xi}{[n+1]}=\bm{\xi}^{*}\}$, such that the process repeats with solving the problem $\mathrm{P}_{3a}^{[i]}[n+2]$. According to part (iii) of Proposition 1 and [sun2016majorization, Theorem 1], every limit point of the sequence $(\mathbf{x}{[0]},\bm{\xi}{[0]}),(\mathbf{x}{[1]},\bm{\xi}{[1]}),\cdots)$ is a stationary point to $\mathrm{P}_{3a}^{[i]}$ as long as the starting point $\{\mathbf{x}{[0]},\bm{\xi}{[0]}\}$ is feasible to $\mathrm{P}_{3a}^{[i]}$.
@@ V-C Algorithm and Complexity
The entire procedure of the PAMM method is summarized in Algorithm 3. In terms of computational complexity, $\mathrm{P}_{3a}^{[i]}[n+1]$ involves $2K$ variables. Therefore, the worst-case complexity for solving $\mathrm{P}_{3a}^{[i]}[n+1]$ is $\mathcal{O}\big((2K)^{3.5}\big)$. Consequently, the total complexity for solving $\mathrm{P}_{3a}^{[i]}$ is $\mathcal{O}\big(\mathcal{J}\,(3K)^{3.5}\big)$, where $\mathcal{J}$ is the number of iterations needed for the $M$ algorithm to converge. On the other hand, since $\mathrm{P}_{3b}^{[i]}[n+1]$ involves $K$ variables, the worst-case complexity for solving $\mathrm{P}_{3b}^{[i]}$ is $\mathcal{O}\big(K^{3.5}\big)$. Based on the above analysis, the total complexity for solving $\mathrm{P}_{3}$ exploiting PAMM is $\mathcal{O}\big(\mathcal{J}\,(3K)^{3.5}+K^{3.5}\big)$.
@@ VI Experiments
We implement the proposed STT-GS system with FDC, PTTM, and PAMM algorithms in Python based on the 3D GS project [kerbl20233d]. The system is deployed on an Ubuntu workstation equipped with a 3.75 GHz AMD EPYC 16-core CPU and an NVIDIA RTX 4090 GPU. We consider the case of $N=64$ and $K=5$, where all clients are randomly distributed within a $100\times 100\,\text{m}^{2}$ area, while the edge server is located at the center $(0,0)$. Then we compute the distances $\{d_{k}\}$ for all $K$ based on the locations of clients and server. The pathloss exponent $\alpha=3$ and the shadow fading $\omega_{k}=-20$ dB. The noise power $\sigma^{2}=-120\,\text{dBm}$. The total time budget is $T=350$ s. The total power is set to $P_{\text{sum}}=300$ mW and the maximum power is $P_{\text{max}}=200$ mW according to [yu2004iterative, zhang2024efficient, schubert2004solution, al2011achieving, beck20141, wang2020learning, wang2020machine, yoo2019learning]. The system bandwidth is set to $B=10\,\text{MHz}$. Simulation parameters are summarized in Table II, where the channel parameters are set according to [wang2020angle].
@@
Experiments are conducted exploiting the rubble-pixsfm LAE dataset [turki2022mega], which consists of $1680$ real-world images. The dataset is divided into 66 batches, where 55 batches are adopted as training data at the 55 clients and the remaining batch is reserved for testing. The data volumes at clients 1–5 are: $2091.26$ MB, $2103.93$ MB, $1906.72$ MB, $1891.08$ MB, and $1544.17$ MB, respectively. We randomly choose one of the 55 batches to train the initial model $\mathcal{S}^{\prime}$ at the edge server.
@@ VI-A1 GS Models
Our implementation builds directly on the open-source 3D GS repository by Kerbl et al. [kerbl20233d], specifically utilizing the 3dgs-accel branch of the diff-gaussian-rasterization submodule for accelerated differentiable rendering.
@@ VI-A3 Initial Model $\mathcal{S}^{\prime}$
The initial model $\mathcal{S}^{\prime}$ was pretrained on the third partition of the rubble-pixsfm dataset used in the current experiments. A sparse 3D point cloud was first reconstructed using COLMAP [schonberger2016structure] via feature extraction, image matching, and triangulation from the RGB images in the selected partition. This SfM-based reconstruction provided the initial 3D keypoints for Gaussian placement. Subsequently, the 3DGS model $\mathcal{S}^{\prime}$ was trained using the official implementation [kerbl20233d], with isotropic Gaussians initialized at the reconstructed keypoints. Training was performed for 40,000 iterations with a spatial resolution factor of 8, under standard settings (Adam optimizer, mixed precision). No client-specific weighting or adaptive sampling was applied during this pretraining phase.
@@ VI-B Evaluation of EGS Loss Prediction
Then, we compare our FDC with two benchmark methods: Equal-interval Sampling and Random Sampling [mayer2020adversarial]. The actual and predicted losses for each client at sampling ratio $\rho=10\%$ are summarized in Table III. The results indicate that the predicted losses of FDC for all clients closely match the ground-truth loss very well. Furthermore, FDC yields the smallest loss discrepancy compared to the other methods. Specifically, the prediction loss error with FDC is within 1.4%, while other methods may involve over 10% discrepancy due to restricted feature diversity of the sampled data. These findings confirm that FDC provides a more reliable and accurate loss prediction.
@@ VI-C Evaluation of EGS Full Transmission
Then, we compare the proposed STT-GS consisting of FDC, PTTM, and PAMM algorithms against the benchmark schemes. We consider the following baselines: 1) MaxRate: EGS with water-filling power allocation for sum-rate maximization [zhang2024efficient]; 2) Fairness: EGS with max-min fairness power allocation [zheng2016wireless]; 3) Active Learning: EGS exploiting only active loss prediction, without considering channel conditions [yoo2019learning].
@@
To gain further insights into the above results, Fig. 11 provides the user selection results of different schemes in a representative scenario. It can be seen that the MaxRate scheme allocates more powers to clients $\{2,3,5\}$, which are the closest users with the most favorable channel conditions. This allocation aligns with its throughput maximization objective, but may result in inefficient use of resources by prioritizing nearby users whose data may be of limited value. The Active Learning scheme selects clients $\{1,5\}$, whose data yields the highest validation losses (thereby greatest information gains) for GS model $\mathcal{E}^{{}^{\prime}}$ as shown in Table III. However, client 11 is situated far from the server, leading to excessive communication costs. By excluding client 11, our proposed method is able to connect more clients, specifically clients $\{2,4,5\}$, all of which posses high-quality GS data as seen in Table III. This corroborates the fact that our method achieves the best tradeoff between GS gains and communication costs, which also accounts for its superior performance in Table III. It is also notable that the Fairness scheme only serves client 55, since client 55 possesses the minimum data volume.
@@ VI-D Sensitivity Analysis
To test the robustness of the proposed method, we train the GS model under different training iterations, i.e., $(7000,10000,20000,30000,40000)$. The trained models are then used to render images, with performance evaluated using PSNR, LPIPS, and SSIM metrics. As can be observed in Fig. 14, the PSNR and SSIM values increase as the number of iterations increases. Specifically, our STT-GS demonstrates superior performance compared to three other methods across all the iterations. We also observe that the rendering performance becomes saturated when the number of iterations exceeds $30,000$. This supports the rationale for a default setting of $30,000$ iterations for GS training, which corroborates the findings in [kerbl20233d].

### v4

@@ IV-A Feature Domain Clustering for Efficient Sampling
Specifically, we consider the Hue Saturation Value (HSV) feature representation [44], where H is hue angle, S is saturation intensity, and V is value brightness. The feature extraction process is thus
@@ V EGS Full Transmission With
Problem $\mathsf{P}_{2}$ involves nonlinear coupling among binary variables $\mathbf{x}$ and continuous variables $\mathbf{p}$. Furthermore, even if we relax $\mathbf{x}$ into a continuous variable, the resultant problem is still nonconvex due to potential multi-user interference. Hence, existing optimization solvers such as Mosek are not applicable. To tackle the challenge, in the following, we will propose a PAMM algorithm that solve $\mathsf{P}_{2}$ efficiently. It consists of an outer penalty alternating minimization (PAM) iteration and an inner majorization minimization (MM) iteration (thus we name it PAMM). Below we present PAM and MM in detail.
@@ V-A Penalty Alternating Minimization
To tackle the discontinuity in constraint (12d), we first relax the binary constraint $x_{k}\in\{0,1\}$ into an affine constraint $x_{k}\in[0,1]$, $\forall k$. However, the relaxation is generally not tight and the solution to the relaxed problem could be $0<x_{k}<1$. To promote a binary solution for the relaxed variable $\{x_{k}\}$, we augment the objective function with a penalty term as in [45]. Here, we adopt the following penalty function [46]:
@@
According to [46, Proposition 1], with the penalty term in (24), there exists an upper bound $\bar{\beta}>0$ such that for any $\beta\in[0,\bar{\beta}]$, $\mathrm{P}_{2}$ and $\mathrm{P_{3}}$ are equivalent with a proper choice of $\beta$ [46].
Now, variables $\{\mathbf{x},\bm{\xi}\}$ and $\mathbf{p}$ are independent in all constraints. Consequently, to solve $\mathrm{P}_{3}$, we can adopt alternating minimization that solves $\{\mathbf{x},\bm{\xi}\}$’s subproblem and $\mathbf{p}$’s subproblem iteratively. This procedure is guaranteed to converge to a local minimum solution of $\mathsf{P}_{3}$ according to [47].
@@ V-B Majorization Minimization
In the iterative procedure (28), $\mathrm{P}_{3b}^{[i]}$ is a linearly constrained quadratic optimization problem and can be optimally solved by off-the-shelf software (e.g., CVXPY). However, $\mathrm{P}_{3a}^{[i]}$ is a nonconvex optimization problem, since the function $\varphi_{1}$ in (34a) and the function $\Phi_{k}$ in (34b) are nonconvex. To tackle these functions, we propose to leverage the framework of MM, which constructs a sequence of upper bounds $\{\widehat{\varphi}_{1}\}$ on $\varphi_{1}(\mathbf{x})$ and replaces $\varphi_{1}(\mathbf{x})$ in $\rm{P}3$ with $\{\widehat{\varphi}_{1}\}$ to obtain the surrogate problems. Similarly, MM would also construct upper bounds $\{\widehat{\Phi}_{k}\}$ on $\Phi_{k}$ and replace $\Phi_{k}$ in $\rm{P}3$ with $\{\widehat{\Phi}_{k}\}$. More specifically, given any feasible solution $\{\mathbf{x}^{\star},\bm{\xi}^{\star}\}$ to $\mathrm{P}3$, we define surrogate functions
@@
Based on part (ii) of Proposition 1, the problem $\mathrm{P}_{3a}^{[i]}[n+1]$ is convex and can be solved by off-the-shelf software packages (e.g., Mosek) for convex programming. Denote its optimal solution as $\{\mathbf{x}^{*},\bm{\xi}\}$. Then we set $\mathbf{x}^{[n+1]}=\mathbf{x}^{*}$ and $\{\bm{\xi}{[n+1]}=\bm{\xi}^{*}\}$, such that the process repeats with solving the problem $\mathrm{P}_{3a}^{[i]}[n+2]$. According to part (iii) of Proposition 1 and [32, Theorem 1], every limit point of the sequence $(\mathbf{x}{[0]},\bm{\xi}{[0]}),(\mathbf{x}{[1]},\bm{\xi}{[1]}),\cdots)$ is a stationary point to $\mathrm{P}_{3a}^{[i]}$ as long as the starting point $\{\mathbf{x}{[0]},\bm{\xi}{[0]}\}$ is feasible to $\mathrm{P}_{3a}^{[i]}$.
@@ V-C Algorithm and Complexity
The entire procedure of the PAMM method is summarized in Algorithm 3. In terms of computational complexity, $\mathrm{P}_{3a}^{[i]}[n+1]$ involves $2K$ variables. Therefore, the worst-case complexity for solving $\mathrm{P}_{3a}^{[i]}[n+1]$ is $\mathcal{O}\big((2K)^{3.5}\big)$. Consequently, the total complexity for solving $\mathrm{P}_{3a}^{[i]}$ is $\mathcal{O}\big(\mathcal{J}\,(3K)^{3.5}\big)$, where $\mathcal{J}$ is the number of iterations needed for the MM algorithm to converge. On the other hand, since $\mathrm{P}_{3b}^{[i]}[n+1]$ involves $K$ variables, the worst-case complexity for solving $\mathrm{P}_{3b}^{[i]}$ is $\mathcal{O}\big(K^{3.5}\big)$. Based on the above analysis, the total complexity for solving $\mathrm{P}_{3}$ exploiting PAMM is $\mathcal{O}\big(\mathcal{J}\,(3K)^{3.5}+K^{3.5}\big)$.
@@ VI Experiments
We implement the proposed STT-GS system with FDC, PTTM, and PAMM algorithms in Python based on the 3D GS project [7]. The system is deployed on an Ubuntu workstation equipped with a 3.75 GHz AMD EPYC 16-core CPU and an NVIDIA RTX 4090 GPU. We consider the case of $N=64$ and $K=5$, where all clients are randomly distributed within a $100\times 100\,\text{m}^{2}$ area, while the edge server is located at the center $(0,0)$. Then we compute the distances $\{d_{k}\}$ for all $K$ based on the locations of clients and server. The pathloss exponent $\alpha=3$ and the shadow fading $\omega_{k}=-20$ dB. The noise power $\sigma^{2}=-120\,\text{dBm}$. The total time budget is $T=350$ s. The total power is set to $P_{\text{sum}}=300$ mW and the maximum power is $P_{\text{max}}=200$ mW according to [10, 11, 12, 13, 14, 15, 16, 17]. The system bandwidth is set to $B=10\,\text{MHz}$. Simulation parameters are summarized in Table II, where the channel parameters are set according to [37].
@@
Experiments are conducted exploiting the rubble-pixsfm LAE dataset [35], which consists of $1680$ real-world images. The dataset is divided into $6$ batches, where $5$ batches are adopted as training data at the $5$ clients and the remaining batch is reserved for testing. The data volumes at clients 1–5 are: $2091.26$ MB, $2103.93$ MB, $1906.72$ MB, $1891.08$ MB, and $1544.17$ MB, respectively. We randomly choose one of the $5$ batches to train the initial model $\mathcal{S}^{\prime}$ at the edge server.
@@ VI-A1 GS Models
Our implementation builds directly on the open-source 3D GS repository by Kerbl et al. [7], specifically utilizing the 3dgs-accel branch of the diff-gaussian-rasterization submodule for accelerated differentiable rendering.
@@ VI-A3 Initial Model $\mathcal{S}^{\prime}$
The initial model $\mathcal{S}^{\prime}$ was pretrained on the third partition of the rubble-pixsfm dataset used in the current experiments. A sparse 3D point cloud was first reconstructed using COLMAP [48] via feature extraction, image matching, and triangulation from the RGB images in the selected partition. This SfM-based reconstruction provided the initial 3D keypoints for Gaussian placement. Subsequently, the 3DGS model $\mathcal{S}^{\prime}$ was trained using the official implementation [7], with isotropic Gaussians initialized at the reconstructed keypoints. Training was performed for 40,000 iterations with a spatial resolution factor of 8, under standard settings (Adam optimizer, mixed precision). No client-specific weighting or adaptive sampling was applied during this pretraining phase.
@@ VI-B Evaluation of EGS Loss Prediction
Then, we compare our FDC with two benchmark methods: Equal-interval Sampling and Random Sampling [43]. The actual and predicted losses for each client at sampling ratio $\rho=10\%$ are summarized in Table III. The results indicate that the predicted losses of FDC for all clients closely match the ground-truth loss very well. Furthermore, FDC yields the smallest loss discrepancy compared to the other methods. Specifically, the prediction loss error with FDC is within 1.4%, while other methods may involve over 10% discrepancy due to restricted feature diversity of the sampled data. These findings confirm that FDC provides a more reliable and accurate loss prediction.
@@ VI-C Evaluation of EGS Full Transmission
Then, we compare the proposed STT-GS consisting of FDC, PTTM, and PAMM algorithms against the benchmark schemes. We consider the following baselines: 1) MaxRate: EGS with water-filling power allocation for sum-rate maximization [11]; 2) Fairness: EGS with max-min fairness power allocation [18]; 3) Active Learning: EGS exploiting only active loss prediction, without considering channel conditions [17].
@@
To gain further insights into the above results, Fig. 11 provides the user selection results of different schemes in a representative scenario. It can be seen that the MaxRate scheme allocates more powers to clients $\{2,3,5\}$, which are the closest users with the most favorable channel conditions. This allocation aligns with its throughput maximization objective, but may result in inefficient use of resources by prioritizing nearby users whose data may be of limited value. The Active Learning scheme selects clients $\{1,5\}$, whose data yields the highest validation losses (thereby greatest information gains) for GS model $\mathcal{E}^{{}^{\prime}}$ as shown in Table III. However, client $1$ is situated far from the server, leading to excessive communication costs. By excluding client $1$, our proposed method is able to connect more clients, specifically clients $\{2,4,5\}$, all of which posses high-quality GS data as seen in Table III. This corroborates the fact that our method achieves the best tradeoff between GS gains and communication costs, which also accounts for its superior performance in Table III. It is also notable that the Fairness scheme only serves client $5$, since client $5$ possesses the minimum data volume.
@@ VI-D Sensitivity Analysis
To test the robustness of the proposed method, we train the GS model under different training iterations, i.e., $(7000,10000,20000,30000,40000)$. The trained models are then used to render images, with performance evaluated using PSNR, LPIPS, and SSIM metrics. As can be observed in Fig. 14, the PSNR and SSIM values increase as the number of iterations increases. Specifically, our STT-GS demonstrates superior performance compared to three other methods across all the iterations. We also observe that the rendering performance becomes saturated when the number of iterations exceeds $30,000$. This supports the rationale for a default setting of $30,000$ iterations for GS training, which corroborates the findings in [7].